OpenStudy (anonymous):

find the values of x & y that maximize or minimize the objective function 4x+3y≥30 x+3y≥21 x≥0,y≥0 Minimum for C=5x=8y

OpenStudy (raden):

graph both lines above : 4x+3y = 30 and x + 3y = 21  to get the coordinat point B, do elimination and subtitution method to solve systems : 4x + 3y = 30 x + 3y = 21 ------------ (-) 3x = 9 x = 9/3 = 3 x + 3y = 21 3 + 3y = 21 3y = 21 - 3 3y = 18 y = 6 so, point B is (3,6)